1.   A population consisting of the values (1, 3, 5, 7). Consider all possible samples of size 2 which can be drawn with replacement from the population. Find the mean and variance of the sampling distribution of the sample mean.


1
Expert's answer
2022-04-25T15:55:46-0400

The number of possible samples which can be selected with replacement is

"N^n=4^2=16."



Population mean:

"\mu=\cfrac{1+3+5+7}{4}=4."


Population variance:

"\sigma^2=\sum(x_i-\mu)^2\cdot P(x_i),"

"X-\mu=\begin{Bmatrix}
 1-4,3-4, 5-4,7-4
\end{Bmatrix}="

"=\begin{Bmatrix}
-3, - 1,1,3
\end{Bmatrix},"

"\sigma^2=(-3)^2\cdot \cfrac{1}{4}+(-1)^2\cdot \cfrac{1}{4}+\\
+1^2\cdot \cfrac{1}{4}+3^2\cdot \cfrac{1}{4}=5."



Mean of the sampling distribution of sample means:

"\mu_{\bar x} =\mu=4."


Variance of the sampling distribution of sample means:

"\sigma^2_{\bar x}=\cfrac{\sigma^2}{n}=\cfrac{5}{2}=2.5."